OpenStudy (anonymous):

An expression is shown below: f(x) = -16x2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. (2 points) Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. (3 points) Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph. (5 points)

OpenStudy (anonymous):

@UsukiDoll this is the last one im having trouble with, can you help me? please

OpenStudy (michele_laino):

part A: we have to solve the subsequent quadratic equation: \[\Large - 16{x^2} + 22x + 3 = 0\]

OpenStudy (anonymous):

can you help me with that? again im new to this...

OpenStudy (michele_laino):

I rewrite that equation as below: \[\Large 16{x^2} - 22x - 3 = 0\]

OpenStudy (michele_laino):

Next I compare that equation with this more general equation: \[\Large a{x^2} + bx + c = 0\] so we can write: a=16, b=-22, and c=-3 am I right?

OpenStudy (anonymous):

yes

OpenStudy (michele_laino):

now, in order to find the requested x-intercepts, you have to apply this formula: \[\Large x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}}\] what do you get?

OpenStudy (anonymous):

completely lost

OpenStudy (michele_laino):

you have to substitute the values of a, b, and c into that formula: \[\Large \begin{gathered} x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}} = \hfill \\ \hfill \\ = \frac{{ - \left( { - 22} \right) \pm \sqrt {{{\left( { - 22} \right)}^2} - 4 \times 16 \times \left( { - 3} \right)} }}{{2 \times 16}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

i dont have a calculator at the moment, its hard doing it in my head

OpenStudy (michele_laino):

please use windows calculator

OpenStudy (michele_laino):

\[\Large \begin{gathered} x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}} = \hfill \\ \hfill \\ = \frac{{ - \left( { - 22} \right) \pm \sqrt {{{\left( { - 22} \right)}^2} - 4 \times 16 \times \left( { - 3} \right)} }}{{2 \times 16}} = \hfill \\ \hfill \\ = \frac{{22 \pm \sqrt {484 + 192} }}{{32}} = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

mine isnt working at the moment...

OpenStudy (michele_laino):

next step: \[\Large \begin{gathered} x = \frac{{ - b \pm \sqrt {{b^2} - 4 \times a \times c} }}{{2 \times a}} = \hfill \\ \hfill \\ = \frac{{ - \left( { - 22} \right) \pm \sqrt {{{\left( { - 22} \right)}^2} - 4 \times 16 \times \left( { - 3} \right)} }}{{2 \times 16}} = \hfill \\ \hfill \\ = \frac{{22 \pm \sqrt {484 + 192} }}{{32}} = \hfill \\ \hfill \\ = \frac{{22 \pm 26}}{{32}} = \begin{array}{*{20}{c}} {\frac{{22 + 26}}{{32}} = ...?} \\ {\frac{{22 - 26}}{{32}} = ...?} \end{array} \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

48/32 and -4/32

OpenStudy (anonymous):

leaves you with -12/32

OpenStudy (michele_laino):

yes! and we can simplify them as below: \[\Large {x_1} = \frac{{48}}{{32}} = \frac{3}{2},\quad {x_2} = - \frac{4}{{32}} = - \frac{1}{8}\]

OpenStudy (michele_laino):

what is -12/32?

OpenStudy (anonymous):

your right, i was overr thinking..

OpenStudy (michele_laino):

so we got your x-intercepts

OpenStudy (michele_laino):

Now, part B: the vertex of a parabola is the point at the top or the point at the bottom of a parabola

OpenStudy (michele_laino):

we have these subsequent cases: